Circular No. 8178

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gx IN MCG +06-3-7 Further to IAUC 8108, T. Boles reports the discovery of an apparent

supernova on unfiltered CCD images taken on Aug. 3.038 (mag 18.6), 4.044 (mag 18.5), and 6.104 UT (limiting mag 20.5). The new object is located at α = 1^{h}02^{m}55^{s}.09, δ = +37^{o}40′10′′.3, which is ≈ 2′′.3 west and 13′′.0 south of the center of MCG +06-3-7. SN 2003gx is not present on Boles’ images from Jan. 16, 23, and Feb. 19 (limiting mag 19.0), and it is not present on Palomar Sky Survey plates from 1989 (blue) and 1991 (red; limiting mag ∼ 20.5).

SUPERNOVAE 2003gy, 2003gz, 2003ha, 2003hb Further to IAUC 8147, P. Astier and C. Pritchet, on behalf of the

CFHT Legacy Survey Supernova Project (plus D. Fouchez and J. Raux), report the discovery of additional apparent supernovae found on CCD images taken during the engineering pre-survey. Additional i′ magnitudes: 2003gy, June 6, [24.5: (S/N = 5 for 2′′ aperture); 26.3, 23.8; 30.3, 23.7; July 4.3, 23.6. 2003gz, June 6, [24.5:; 26.3, 22.8; 30.3, 23.0; July 4.3, 23.0. 2003ha, June 5, [24.5:; 26.5, 22.0; 30.5, 21.5, July 4.5, 21.2. 2003hb, June 5, [24.5:; 26.5, 21.6; 30.5, 21.8, July 4.5, 21.9.

SN  2003 UT  α_{2000}  δ_{2000}  i′  Offset
2003gy June 22.3 14^{h}21^{m}27.^{s}76 +53^{o}06′10′′.3 24.0 0′′.6 E, 1′′.1 N
2003gz June 22.3 14 22 39.87 +52 47 49.5 23.6 0′′.2 W, 0′′.3 S
2003ha June 22.5 22 14 45.80 −17 44 22.9 22.9 0′′.7 W, 0′′.6 S
2003hb June 22.5 22 16 56.31 −17 57 37.7 21.8 0′′.7 W, 0′′.2 N

COMETS C/2003 H5, C/2003 H6, C/2003 H7 (SOHO) Further to IAUC 8138, D. Hammer has reported astrometric measure-

ments for several additional comets found on C2 coronagraph images at the SOHO website by R. Kracht. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2003-P02. C/2003 H6 and C/2003 H7 are obviously related to each other; C/2003 H5 belongs to the Meyer group.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 H5 Apr. 27.268  2^{h}22.^{m}6  +14^{o}49′
C/2003 H6  30.201  2 24.3  +14 23
C/2003 H7  30.129  2 24.8  +14 16
2003 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green